a cyclist comes to a stop from a velocity of 15 m/s the acceleration is -3m/s squared by 2 how far does the bicycle travel while it is braking


1
Expert's answer
2021-02-10T10:07:13-0500
"v_f^2=v_i^2+2ad,""0=v_i^2+2ad""d=\dfrac{-v_i^2}{2a}=\dfrac{-(15\ \dfrac{m}{s^2})^2}{2\cdot(-3\ \dfrac{m}{s^2})}=37.5\ m."